OpenStudy (vishweshshrimali5):

Find all positive integers a, b for which the number (√2+√a)/(√3+√b) is a rational number

OpenStudy (vishweshshrimali5):

@ganeshie8 @ikram002p

OpenStudy (vishweshshrimali5):

Let, m = (√a + √2)/(√b + √3) - (1) => m = (√a + √2)(√b - √3)/(b - 3); if b - 3 is non zero => m = {√(ab) - √(3a) + √(2b) - √(6)}/(b-3) - (2) Now, for m to be a rational number: (i) √6 must cancel out. (ii) No square roots must be present. Lets first see the (i) point: For √6 to cancel out, one of √2b and √(ab) must be equal to √6 Thus, √(2b) = √6 or √(ab) = √6 Case - 1: √(2b) = √6 => 2b = 6 => b = 3 - (3) Substituing the value of b in equation (1). m = (√a + √2)/(2√3) => m = (√(3a) + √6)/6 - (4) Now, m can be a rational number only if √(3a) + √(6) is a rational number, which is not possible for any value of a. Thus, if b = 3, then there is no solution to equation (1). Case - 2: √(ab) = √(6) => ab = 6 Since a and b are integers and must be non negative. Thus, possible solutions to above equation are: (a,b) = (1,6); (2,3); (3,2); (6,1) Now, we check each solution by substituting the values of a and b in equation (1). Only (3,2) and (6,1) satisfy the equation. It was what I attempted. I realise that there is a mistake in this thing: "For √6 to cancel out, one of √2b and √(ab) must be equal to √6" Any idea about how to solve it ?

OpenStudy (vishweshshrimali5):

@hartnn

OpenStudy (vishweshshrimali5):

@Kainui

OpenStudy (paxpolaris):

\[ \sqrt a + \sqrt2 \over \sqrt b + \sqrt 3\] rationalize the denominatior\[={\left( \sqrt a+ \sqrt2 \right)\left( \sqrt b-\sqrt 3\right) \over b-3}\]

OpenStudy (vishweshshrimali5):

Ok

ganeshie8 (ganeshie8):

*

OpenStudy (paxpolaris):

now looking only at numertor:\[=\sqrt{ab}+\sqrt{2b}-\sqrt{3a}-\sqrt6\]

OpenStudy (vishweshshrimali5):

Okay.

OpenStudy (paxpolaris):

a=3, b=2 works

OpenStudy (paxpolaris):

b can't be 3

OpenStudy (vishweshshrimali5):

Yes

OpenStudy (vishweshshrimali5):

I got that. But, what I can't get is that whether (3,2) and (6,1) are the only possible solutions or are there any more ?

OpenStudy (paxpolaris):

=the only other solution could be \[\sqrt{2b}-\sqrt{3a}=\sqrt6\] and ab is perfect square... both are not possible at the same time. so you have all the answers

OpenStudy (vishweshshrimali5):

Okay. Thanks a lot @PaxPolaris :)

OpenStudy (ikram002p):

:O nice question :)